Identify the region/area under the normal curve corresponding to each of the following.


1.Between Z=0 and z=1.36


2.between z=0 and z=1.87


3.between z=1.36 and z= 2.5


4.between -1.36 and z= -1.87


5. Between z=-1.36 and z=2.5


6.between z=1.36 and z=-2.5

1
Expert's answer
2022-04-17T10:12:14-0400

We find from z-table:

"1. \ P(0<Z<1.36) =P(Z<1.36)-P(Z<0)=\\
=0.9131-0.5=0.4131;"

"2. \ P(0<Z<1.87) =P(Z<1.87)-P(Z<0)=\\
=0.9693-0.5=0.4693;"

"3. \ P(1.36<Z<2.5) =\\
=P(Z<2.5)-P(Z<1.36)=\\
=0.9938-0.9131=0.0807;"

"4. \ P(-1.87<Z<-1.36) =\\
=P(Z<-1.36)-P(Z<-1.87)=\\
=0.0869-0.0307=0.0562;"


"5. \ P(-1.36<Z<2.5) =\\
=P(Z<2.5)-P(Z<-1.36)=\\
=0.9938-0.0869=0.9069;"


"6. \ P(-2.5<Z<1.36) =\\
=P(Z<1.36)-P(Z<-2.5)=\\
=0.9131-0.0062=0.9069."